# Vehicle Rebalancing Under Adherence Uncertainty

### v1

@@ II-C1 Supply Vector
$\mathbf{X_{cj}}=\begin{cases}1~{}&{\text{with probability}\hskip 2.84544pt\mathbb{P}_{c}(\{\omega_{2}=j\})}\\ 0~{}&{\text{with probability}\hskip 2.84544pt1-\mathbb{P}_{c}(\{\omega_{2}=j\})}\end{cases}$ (10)
@@
$\mathbb{P}_{c}(\{\omega_{2}=j\})$ $=\mu(c)\cdot x_{cj}+(1-\mu(c))\cdot\mathcal{L}_{cj}$ (11)
@@
$\mathbf{s_{j}}=\sum_{c\in\mathcal{C}}\mathbf{X_{cj}}$ (12)
@@
$\sum_{\begin{subarray}{c}C_{b}\subseteq\mathcal{C}\end{subarray}}\bigg{(}\prod_{c\in C_{b}}\mathbb{P}_{c}(\{\omega_{2}=j\})\prod_{c\notin C_{b}}(1-\mathbb{P}_{c}(\{\omega_{2}=j\}))\bigg{)}$ (13)
@@
$\mathbb{E}[\mathbf{s_{j}}]$ $=\sum_{c=1}^{m}\mathbb{E}[\mathbf{X_{cj}}]=\sum_{c\in C}\mu(c)\cdot x_{cj}+(1-\mu(c))\cdot\mathcal{L}_{cj}$ (14)
@@ II-C2 Demand Vector
Demand vector $\mathbf{d}\in\mathbb{Z}^{|\mathcal{R}|}$ is a random vector consisting of the number of requests that will appear in each region in the planning horizon. Demand is assumed to be a Gaussian random variable, whose parameters are unknown and need to be estimated.
$\mathbf{d_{j}}\sim\mathcal{N}\left(\nu_{j},\sigma_{j}^{2}\right)=\underbrace{\nu_{j}}_{\text{forecast}}+\underbrace{\mathcal{N}(0,\sigma_{j}^{2})}_{\text{error}}$ (15)
First, a function $g_{j}:\mathbf{\tilde{d}_{j}}\to\nu_{j}$ is learned using a gradient-boosted tree-based regression model [26] for each region $j\in\mathcal{R}$ using the historical demand data [23]. This function maps a vector $\mathbf{\tilde{d}_{j}}$, that contains tuples of demand and covariate information of previous planning horizons to the forecast:
$g_{j}(\mathbf{\tilde{d}_{j}})=\nu_{j}$ (16)
where,
@@
Each variable $d_{j}(h-t)$ denotes the realization of the random variable $\mathbf{d_{j}(h-t)}$, i.e., the number of passenger demands that appeared in the region $j\in\mathcal{R}$ in the planning horizon $h-t$ and $\hat{d}(h-t)$ denotes the covariate information, which contains three additional features, namely i) hour of the day, ii) day of the week, and iii) day of the month. Walk-forward validation [27] is used to refine the trained model as the new data corresponding to the current time step is available.
Parameter $\sigma_{j}^{2}$ of the error distribution $\mathcal{N}(0,\sigma_{j}^{2})$ is estimated using the residuals from the test dataset, and is defined as the expected value of the squared residuals:
$\sigma_{j}^{2}=\mathbb{E}[(\mathbf{d_{j}}-\nu_{j})^{2}]$ (17)
@@ II-C3 Travel Time Matrix
Travel time matrix $\mathbf{T}\in\mathbb{R}^{n\times n}$ is a random matrix, where $\mathbf{T_{ij}}$ is the time it will take to travel from region $i\in\mathcal{R}$ to region $j\in\mathcal{R}$. $\mathbf{T_{ij}}$ is modelled as a Gaussian ransom variable:
$\mathbf{T_{ij}}\sim\mathcal{N}(\tau_{ij},\varepsilon_{ij}^{2})$ (18)
@@
$x_{cj}\cdot\big{(}\mathcal{T}_{cj}-|\mathcal{H}|\big{)}\leq 0$ (22)
Further, to avoid over-saturation, i.e., dispatching more taxis than that are required, we use the approach similar to [7], and define an upper limit of the total number of repositioning recommendations.
$\sum_{c\in\mathcal{C}}x_{cj}\leq\rho\cdot\mathbb{E}[\mathbf{d_{j}}]$ (23)
@@ II-C4 Objective Function
$\mathbb{E}\big{[}\min\big{(}\mathbf{s_{j}},\mathbf{d_{j}}\big{)}\big{]}$ (24)
where $\mathbf{s_{j}}$ is a Poisson binomial distributed random variable and $\mathbf{d_{j}}$ is a normally distributed random variable as discussed previously. Obtaining a closed-form expression for $\mathbb{E}\big{[}\min\big{(}\mathbf{s_{j}},\mathbf{d_{j}}\big{)}\big{]}$ is practically intractable due to the computational complexity of deriving the probability distribution of $\mathbf{s_{j}}$. Therefore, we rely on the expected supply $\mathbb{E}[\mathbf{s_{j}}]$, and define an approximation to the expected allocation. This is determined using the expected supply $\mathbb{E}[\mathbf{s_{j}}]$ and expected demand $\mathbb{E}[\mathbf{d_{j}}]$:
@@
The objective is to maximize the expected revenue, i.e., the anticipated earnings from successful allocations. The expected earnings is defined as the product of expected number of allocations $\xi_{j}$ and the expected ride fare $\mathcal{P}_{j}$ (calculated using  [23]) in each region and is given by:
$\sum_{j\in\mathcal{R}}\xi_{j}\cdot\mathcal{P}_{j}$ (26)
The cost $\mathcal{Q}_{cj}$ of moving a taxi from its current location, say region $i\in\mathcal{R}$ to region $j\in\mathcal{R}$, is calculated similarly to  [8], based on the distance between the regions, the average taxi mileage, and the fuel cost. For a driver, the profit is defined as the difference between the earning and costs. Note that the cost of operation is not considered in the objective function, this is because the taxi companies operating in the Uber style model do not bear the transportation cost, rather it the driver who has to bear this cost. The company gets a fixed share of the earnings, and from the companies point of view maximizing overall earnings is a desirable objective.
II-C5 Repositioning Recommendation Model
The repositioning recommendation model (27) generates an optimal taxi-destination match, ensuring the taxis are dispatched to destinations with a higher chance of getting allocated during the planning horizon. Using Eq. 6, Eq. 22, and Eq. 23 as the constraints and expected profit maximization as the objective function, where $\xi_{j}$ is given by Eq. 25 the proposed model can be written as:
@@ II-C5 Repositioning Recommendation Model
Optimal Taxi Rebalancing Recommendation (OTRR).
$\max_{x}$ $\sum_{j\in\mathcal{R}}\min\big{(}\mathbb{E}[\mathbf{s_{j}}],\mathbb{E}[\mathbf{d_{j}}]\big{)}\cdot\mathcal{P}_{j}$ (27)
s.t. $\sum_{j\in\mathcal{R}}x_{cj}=1\hskip 2.84544pt\forall c\in\mathcal{C}$
$\sum_{c\in\mathcal{C}}x_{cj}\leq\rho\cdot\mathbb{E}[\mathbf{d_{j}}]\hskip 2.84544pt\forall j\in\mathcal{R}$
$x_{cj}\cdot\big{(}\mathcal{T}_{cj}-|\mathcal{H}|\big{)}\leq 0\hskip 2.84544pt\forall c\in\mathcal{C},j\in\mathcal{R}$
$x_{cj}\in\{0,1\}\hskip 2.84544pt\forall c\in\mathcal{C},j\in\mathcal{R}.$
The proposed model can be converted to a linear program by using an auxiliary variable $\mathcal{Z}_{j}$ such that, the objective function now becomes the product of $\mathcal{Z}_{j}$ and $\mathcal{P}_{j}$ and two new constraints are introduced:
$\mathcal{Z}_{j}\leq\mathbb{E}[d_{j}]\hskip 2.84544pt\forall j\in\mathcal{R}$ (28)
$\mathcal{Z}_{j}\leq\mathbb{E}[s_{j}]\hskip 2.84544pt\forall j\in\mathcal{R}.$
Note that $\mathbb{E}[d_{j}]=\nu_{j}$ is the output of the demand prediction model as defined by Eq. 16, and $\mathbb{E}[s_{j}]$ is a function of driver preference probability $\mathcal{L}_{cj}$ Eq. 3, driver confidence level Eq. 9, and decision variable $x_{cj}$ Eq. 5 as defined by Eq. 14. Additionally, the constraint Eq. 6 can be relaxed with an inequality and decision variable $x_{cj}$ can also be relaxed to $x_{cj}\in[0,1]$. Incorporating the above mentioned relaxations, the model Eq. 27 can be converted to:
Problem II.2.
Linearized OTRR:
$\max_{x,\mathcal{Z}}$ $\sum_{j\in\mathcal{R}}\mathcal{Z}_{j}\cdot\mathcal{P}_{j}$ (29)
s.t. $\sum_{j\in\mathcal{R}}x_{cj}\leq 1\hskip 2.84544pt\forall c\in\mathcal{C}$
$\sum_{c\in\mathcal{C}}x_{cj}\leq\rho\cdot\mathbb{E}[\mathbf{d_{j}}]\hskip 2.84544pt\forall j\in\mathcal{R}$
$x_{cj}\cdot\big{(}\mathcal{T}_{cj}-|\mathcal{H}|\big{)}\leq 0\hskip 2.84544pt\forall c\in\mathcal{C},j\in\mathcal{R}$
$\mathcal{Z}_{j}\leq\nu_{j}\hskip 2.84544pt\forall j\in\mathcal{R}$
$\mathcal{Z}_{j}\leq\sum_{c\in C}\mu(c)\cdot x_{cj}+(1-\mu(c))\cdot\mathcal{L}_{cj}\hskip 2.84544pt\forall j\in\mathcal{R}$
$x_{cj}\in[0,1]\hskip 2.84544pt\forall c\in\mathcal{C},j\in\mathcal{R}$
$\mathcal{Z}_{j}\in\mathbb{R}_{+}\hskip 2.84544pt\forall j\in\mathcal{R}.$
III Algorithm
The proposed algorithm takes the following inputs, i) confidence parameters for each driver: $\alpha_{r}(c),\beta_{r}(c)$, $\alpha_{p}(c),\beta_{p}(c)$, ii) a preference prediction model for each driver: $\sigma_{c}$, iii) features associated with each region for preference prediction: $\mathbf{z_{j}}$, iv) features associated with each region for demand prediction: $\mathbf{\tilde{d_{j}}}$, v) a demand prediction model of each region:$g_{j}$, vi) features associated with each region pair for travel time prediction $\mathbf{\tilde{T}_{ij}}$, vii) a travel time prediction model: h, and viii) expected profits associated with each region $j\in\mathcal{R}$ in planning horizon: $\mathcal{P}_{j}$; and returns the repositioning recommendations $x_{cj}$ for each idle-standing driver. This algorithm is executed every $\mathcal{H}$ mins. at the beginning the planning horizon.
Algorithm 1 Driver State Aware Reposition Recommendation
1:  Input: $\{\alpha_{r}(c),\beta_{r}(c),\alpha_{p}(c),\beta_{p}(c)\forall c\},\{\sigma_{c}\forall c\},\{\mathbf{z_{j}}\forall j\},$
2:  $\{\mathbf{\tilde{d}_{j}}\forall j\},\{g_{j}\forall j\},\{\mathbf{\tilde{T}_{ij}}\forall i,j\},h,\{\mathcal{P}_{j}\forall j\},$
3:  Output: $x$
4:  Evaluate preference of idle drivers using Eq. 3
5:  for $c\in\mathcal{C}$ do
@@ III Algorithm
7:        $\mathcal{L}_{cj}\leftarrow\text{get\_driver\_preference}(\sigma_{c},\mathbf{z_{j}})$
@@
10:  Evaluate confidence-level of idle drivers using Eq. 9
11:  for $c\in\mathcal{C}$ do
12:     $\mu(c)\leftarrow\text{get\_driver\_confidence}(\alpha_{r}(c),\beta_{r}(c),\alpha_{p}(c),\beta_{p}(c))$
@@
14:  Evaluate expected demand in each region using Eq. 16
15:  for $j\in\mathcal{R}$ do
16:     $\nu_{j}\leftarrow\text{get\_expected\_demand}(g_{j},\mathbf{\tilde{d}})$
17:  end for
18:  Evaluate expected travel time using Eq. 19
19:  for $i\in\mathcal{R}$ do
20:     for $j\in\mathcal{R}$ do
21:        $\tau_{ij}\leftarrow\text{get\_expected\_travel\_time}(h,\mathbf{\tilde{T}_{ij}})$
22:     end for
23:  end for
24:  Evaluate travel time for drivers using Eq. 21
25:  for $c\in\mathcal{C}$ do
26:     $i\leftarrow\text{get\_current\_region\_of\_driver}(c)$
27:     for $j\in\mathcal{R}$ do
28:        $\mathcal{T}_{cj}\leftarrow\text{get\_travel\_time\_for\_driver}(\tau_{ij})$
29:     end for
@@
31:  Evaluate repositioning recommendations by solving (29)
32:  $x\leftarrow\text{get\_recommendations}(\mathcal{L},\mathbf{\mu},\mathbf{\nu},\mathcal{T},\mathcal{P})$
33:  for $c\in\mathcal{C}$ do
34:     if $y_{c}=1$ then
35:        $\alpha_{r}(c)\leftarrow\alpha_{r}(c)+\epsilon_{1}$
36:     else
37:        $\beta_{r}(c)\leftarrow\beta_{r}(c)+\epsilon_{0}$
38:     end if
39:  end for
40:  Evaluate Metric Eq. 30, Eq. 31, Eq. 32, Eq. 33
41:  Return: $x$

### v2

@@ II-C1 Supply Vector
$\mathbf{X_{cj}}=\begin{cases}1~&{\text{with probability}\hskip 2.84544pt\mathbb{P}_{c}(\{\omega_{2}=j\})}\\ 0~&{\text{with probability}\hskip 2.84544pt1-\mathbb{P}_{c}(\{\omega_{2}=j\})}.\end{cases}$ (10)
@@
$\mathbb{P}_{c}(\{\omega_{2}=j\})$ $=\mu(c)\cdot\mathbf{x_{cj}}+(1-\mu(c))\cdot L_{cj}.$ (11)
@@
$\mathbf{s_{j}}=\sum_{c\in\mathcal{C}}\mathbf{X_{cj}}.$ (12)
@@
$\sum_{\begin{subarray}{c}C_{b}\subseteq\mathcal{C}\end{subarray}}\bigg{(}\prod_{c\in C_{b}}\mathbb{P}_{c}(\{\omega_{2}=j\})\prod_{c\notin C_{b}}(1-\mathbb{P}_{c}(\{\omega_{2}=j\}))\bigg{)},$ (13)
@@
$\mathbb{E}[\mathbf{s_{j}}]$ $=\sum_{c=1}^{m}\mathbb{E}[\mathbf{X_{cj}}]=\sum_{c\in C}\mu(c)\cdot\mathbf{x_{cj}}+(1-\mu(c))\cdot L_{cj}.$ (14)
@@ II-C2 Demand Vector
We model the demand vector $\mathbf{d}\in\mathbb{Z}^{|\mathcal{R}|}$ as a random vector representing the number of passenger requests in each region during the planning horizon. Instead of modeling the event-level arrival process (e.g., via a Poisson process), we adopt a forecast-based approach: the expected demand $\nu_{j}$ in each region $j$ is predicted using a supervised regression model (XGBoost), and the uncertainty in this forecast is captured via a Gaussian error model.
Formally, the demand is modeled as:
$\mathbf{d_{j}}\sim\mathcal{N}\left(\nu_{j},\sigma_{j}^{2}\right)=\underbrace{\nu_{j}}_{\text{forecast}}+\underbrace{\mathcal{N}(0,\sigma_{j}^{2})}_{\text{error}}.$ (15)
This formulation reflects a common assumption in forecast-driven optimization, where the prediction error is treated as zero-mean Gaussian noise particularly when probabilistic forecasts are not directly available from the model.
The forecast $\nu_{j}$ is generated by a function $g_{j}:\mathbf{\tilde{d}_{j}}\to\nu_{j}$, trained on historical demand and covariates for each region $j$:
$g_{j}(\mathbf{\tilde{d}_{j}})=\nu_{j},$ (16)
where
@@
. and $\hat{d}(h-t)$ includes temporal features: hour of the day, day of the week, and day of the month. The variance $\sigma_{j}^{2}$ of the forecast error is estimated using residuals from the test set:
$\sigma_{j}^{2}=\mathbb{E}[(\mathbf{d_{j}}-\nu_{j})^{2}].$ (17)
@@ II-C3 Travel Time Matrix
Travel time matrix $\mathbf{T}\in\mathbb{R}^{n\times n}$ is a random matrix, where $\mathbf{T_{ij}}$ is the time it will take to travel from region $i\in\mathcal{R}$ to region $j\in\mathcal{R}$. $\mathbf{T_{ij}}$ is modelled as a Gaussian random variable:
$\mathbf{T_{ij}}\sim\mathcal{N}(\tau_{ij},\varepsilon_{ij}^{2}).$ (18)
@@
$\mathbf{x_{cj}}\cdot\big{(}\mathcal{T}_{cj}-\mathcal{H}\big{)}\leq 0.$ (22)
@@ II-C4 Objective Function
$\mathbb{E}\big{[}\min\big{(}\mathbf{s_{j}},\mathbf{d_{j}}\big{)}\big{]}.$ (23)
where $\mathbf{s_{j}}$ is a Poisson binomial random variable and $\mathbf{d_{j}}$ is a normal random variable as discussed previously. Obtaining a closed-form expression for $\mathbb{E}\big{[}\min\big{(}\mathbf{s_{j}},\mathbf{d_{j}}\big{)}\big{]}$ is practically intractable due to the computational complexity of deriving the probability distribution of $\mathbf{s_{j}}$. Therefore, we rely on the expected supply $\mathbb{E}[\mathbf{s_{j}}]$, and define an approximation to the expected allocation. This is determined using the expected supply $\mathbb{E}[\mathbf{s_{j}}]$ and expected demand $\mathbb{E}[\mathbf{d_{j}}]$:
@@
The first of the two objectives in the proposed model is to maximize the total expected allocations across the network. This is desirable from the ride-hailing platform’s perspective because higher expected allocations translate to higher earnings for the platform. We use the following expression to define the total expected allocations:
$J_{1}=\sum_{j\in\mathcal{R}}\min\big{(}\mathbb{E}[\mathbf{s_{j}}],\mathbb{E}[\mathbf{d_{j}}]\big{)}.$ (25)
Additionally, the ride-hailing platforms also care about the passenger waiting time. Although increasing taxi availability which translates to higher allocation also improves the expected waiting time for the passengers, it is possible to achieve the same level of supply distributions by executing different set of recommendations. This requires explicitly ensuring that the drivers closest to the potential pick-up hot spots must be preferred over the drivers that need to travel for longer periods of time. This is achieved by minimizing the total idle cruising time before allocation by the drivers who have been recommended to reposition using the following expression:
$J_{2}=-\sum_{c\in\mathcal{C}}\sum_{j\in\mathcal{R}}\mathbf{x_{cj}}\cdot\mathcal{T}_{cj}.$ (26)
II-C5 Optimization Model
The proposed model generates an optimal taxi-destination match, ensuring the taxis are dispatched to destinations with a higher chance of getting allocated during the planning horizon while ensuring the traversal time to reach the destination is also minimized. Using Eq. 6, Eq. 22, as the constraints and Eq. 25 and Eq. 26 as the objective functions, the proposed Adherence Aware Vehicle Rebalance Model (AAVR) is defined as follows:
@@ II-C5 Optimization Model
Adherence Aware Vehicle Rebalancing Model (AAVR).
$\max_{\mathbf{x}\in\{0,1\}^{|\mathcal{C}|\times|\mathcal{R}|}}$ $\sum_{j\in\mathcal{R}}\min\big{(}\mathbb{E}[\mathbf{s_{j}}],\mathbb{E}[\mathbf{d_{j}}]\big{)}-\beta\sum_{j\in\mathcal{R}}\sum_{c\in\mathcal{C}}\mathbf{x_{cj}}\cdot\mathcal{T}_{cj}$ (27)
s.t. $\sum_{j\in\mathcal{R}}\mathbf{x_{cj}}=1\hskip 2.84544pt\forall c\in\mathcal{C}$
$\mathbf{x_{cj}}\cdot\big{(}\mathcal{T}_{cj}-\mathcal{H}\big{)}\leq 0\hskip 2.84544pt\forall c\in\mathcal{C},j\in\mathcal{R}$
The above problem can be converted into a MILP as discussed in Problem IV.1.
III AAVR - Model Architecture and Algorithm
Figure 3: Architecture of the Adherence-Aware Vehicle Rebalancing (AAVR) model. The system comprises four components: (i) an adherence estimator, combining a Beta–Bernoulli bandit with Thompson Sampling and a multinomial logit model, to compute driver-specific acceptance probabilities; (ii) a supply–demand estimator that forecasts adherence-adjusted supply and future demand using region-level XGBoost models; (iii) a travel-time estimator that predicts inter-region travel times via a network-level XGBoost model; and (iv) an optimization module that selects repositioning actions by maximizing expected allocations and minimizing idle travel time.
Figure 3 presents the architecture of the proposed AAVR model, which consists of four core modules: adherence estimation, supply-demand aggregation, travel time prediction, and optimization. The first module estimates the likelihood that each idle-standing driver $\mathrm{DRV}_{i}$ will accept a repositioning recommendation. This adherence probability, is computed by combining two driver-specific models: a Beta-Bernoulli Bandit with Thompson Sampling (BBB-TS), which captures the driver’s confidence in the system’s recommendations based on historical acceptance data (Eq. 9), and a Logit model that estimates the driver’s spatial preferences over regions using regional features (Eq. 3). These outputs are combined to form the adherence probability conditioned on the recommended region (Eq. 11). The second module aggregates supply and demand information across regions. The expected supply in a region is calculated using the number of idle drivers and their adherence probabilities (Eq. 14), while demand is predicted using an XGBoost model trained on historical demand data for each region. The expected number of successful allocations is then derived from these estimates (Eq. 24), and the sum over all regions constitutes the first objective of the optimization problem: maximizing expected total allocations (Eq. 25). The third module estimates inter-region travel times using a network-level XGBoost model trained on historical trip durations. These estimates are used to calculate the time required for each driver to reach any recommended region (Eq. 21), contributing to both the second objective minimizing total idle travel time (Eq. 26) and a feasibility constraint that ensures drivers can reach their assigned regions within the planning horizon (Eq. 22). The final module solves the optimization problem defined in Problem II.1, which balances the trade-off between maximizing expected allocations and minimizing idle travel time while respecting supply, demand, and reachability constraints.
The AAVR - Algorithm 1 is executed at the beginning of each planning horizon, every $\mathcal{H}$ minutes, and takes as input: (i) BBB-TS (Beta Bernoulli Bandit with Thompson Sampling) model for each driver $c\in\mathcal{C}$; (ii) Logit model $\sigma_{c}$ for each driver $c\in\mathcal{C}$; (iii) feature vector $\mathbf{z}_{j}$ for preference prediction and $\mathbf{\tilde{d}}_{j}$ for demand forecasting for each region $j\in\mathcal{R}$; (iv) a demand prediction model $g_{j}$ for each region $j\in\mathcal{R}$; (v) pairwise region features $\mathbf{\tilde{T}}_{ij}$ and a travel time prediction model $h$. Using these inputs, the algorithm produces binary repositioning decisions $\mathbf{x_{cj}}$, indicating whether driver $c$ should reposition to region $j$, with the goal of improving system-level efficiency while accounting for individual driver behavior and spatial preferences. Upon execution, performance metrics such as the total allocation, average wait time, platform profit, and driver profits are calculated to evaluate recommendations.
The overall computational complexity of the AAVR algorithm arises from five main components. First, inter-region travel time estimation incurs a cost of $\mathcal{O}(|\mathcal{R}|^{2})$ due to pairwise computations between all region pairs. Second, demand forecasting contributes $\mathcal{O}(|\mathcal{R}|)$ operations using region-level models. Third, preference evaluation and driver-specific travel time computation require $\mathcal{O}(|\mathcal{C}|\cdot|\mathcal{R}|)$ operations to compute compatibility scores and travel estimates for each driver-region pair. Fourth, the repositioning optimization is formulated as a mixed-integer linear program (MILP) with $|\mathcal{C}|\cdot|\mathcal{R}|$ binary decision variables. While MILPs are NP-hard in the worst case, modern solvers can handle instances of practical size efficiently using branch-and-bound-based algorithms. Finally, updating the confidence model parameters incurs an additional $\mathcal{O}(|\mathcal{C}|)$ cost.
Algorithm 1 Adherence-Aware Vehicle Rebalancing
[…]
2:Output: Rebalancing decision matrix $\mathbf{x}\in\{0,1\}^{|\mathcal{C}|\times|\mathcal{R}|}$
3:Preference & Confidence Estimation
4:for $c\in\mathcal{C}$ do
[…]
@@ III AAVR - Model Architecture and Algorithm
7:   $L_{cj}\leftarrow\text{get\_driver\_preference}(\sigma_{c},\mathbf{z}_{j})$
@@
10:Demand & Travel Time Estimation
11:for $j\in\mathcal{R}$ do
12:  $\nu_{j}\leftarrow\text{get\_expected\_demand}(g_{j},\mathbf{\tilde{d}}_{j})$
@@
14:for $c\in\mathcal{C}$ do
15:  $i\leftarrow\text{get\_current\_region}(c)$
16:  for $j\in\mathcal{R}$ do
17:   $\tau_{ij}\leftarrow\text{get\_expected\_travel\_time}(h,\mathbf{\tilde{T}}_{ij})$
18:   $\mathcal{T}_{cj}\leftarrow\tau_{ij}$
19:  end for
20:end for
21:Solve Optimization Problem (Eq. 28)
[…]
23:Bandit Update
24:for $c\in\mathcal{C}$ do
25:  if $y_{c}=1$ then
26:   $\alpha_{r}(c)\leftarrow\alpha_{r}(c)+\epsilon_{1}$
27:  else
28:   $\beta_{r}(c)\leftarrow\beta_{r}(c)+\epsilon_{0}$
29:  end if
@@
31:Return: $\mathbf{x}$